# EEG-MACS: Manifold Attention and Confidence Stratification for EEG-based Cross-Center Brain Disease Diagnosis under Unreliable Annotations

### v1

@@ Abstract.
Cross-center data heterogeneity and annotation unreliability significantly challenge the intelligent diagnosis of diseases using brain signals. A notable example is the EEG-based diagnosis of neurodegenerative diseases, which features subtler abnormal neural dynamics typically observed in small-group settings. To advance this area, in this work, we introduce a transferable framework employing Manifold Attention and Confidence Stratification (MACS) to diagnose neurodegenerative disorders based on EEG signals sourced from four centers with unreliable annotations. The MACS framework’s effectiveness stems from these features: 1) The Augmentor generates various EEG-represented brain variants to enrich the data space; 2) The Switcher enhances the feature space for trusted samples and reduces overfitting on incorrectly labeled samples; 3) The Encoder uses the Riemannian manifold and Euclidean metrics to capture spatiotemporal variations and dynamic synchronization in EEG; 4) The Projector, equipped with dual heads, monitors consistency across multiple brain variants and ensures diagnostic accuracy; 5) The Stratifier adaptively stratifies learned samples by confidence levels throughout the training process; 6) Forward and backpropagation in MACS are constrained by confidence stratification to stabilize the learning system amid unreliable annotations. Our subject-independent experiments, conducted on both neurocognitive and movement disorders using cross-center corpora, have demonstrated superior performance compared to existing related algorithms. This work not only improves EEG-based diagnostics for cross-center and small-setting brain diseases but also offers insights into extending MACS techniques to other data analyses, tackling data heterogeneity and annotation unreliability in multimedia and multimodal content understanding. We have released our code here: https://anonymous.4open.science/r/EEG-Disease-MACS-0B4A.
@@ 1. Introduction
Addressing the challenges posed by data heterogeneity and annotation unreliability is essential for the analysis and interpretation of multimedia and multimodal data [1, 2]. These issues are particularly prevalent in human-centered computing, where they encounter even greater difficulties in small-group settings. A prime example is the analysis of neural dynamics, such as the automated diagnosis of cognitive and movement brain disorders through electroencephalogram (EEG) signals  [3, 4]. While EEG-based diagnosis models have undergone extensive research, they are often disease-specific, rely on single-center studies, and depend on fully supervised learning with annotated data  [5]. Developing effective and transferable models capable of handling cross-center data and low-quality annotations is crucial for enhancing adaptability across different diseases and advancing EEG modeling techniques.
@@
Current EEG analyzing methods range from traditional feature engineering to deep learning approaches. Various features extracted from domains like time (e.g., Hjorth parameters), frequency (e.g., Fourier Transform), dynamics (e.g., entropy), and functional networks (e.g., synchronization and phase coupling) have demonstrated efficacy due to EEG’s intrinsic characteristics [6, 7, 8, 9]. The sparsity of handcrafted features has led to the adoption of deep learning models, such as convolutional neural networks, graph neural networks, and hybrid approaches, to capture spatial-temporal patterns, structured characteristics, and other high-level representations [10, 11, 12, 13]. However, these existing methods often struggle with noisy data and lack cross-center applicability. This may primarily be due to: 1) the ineffective mapping that cannot project EEG signals into a high-level feature space invariant across diverse data distributions from multiple centers; 2) the absence of integrating advanced learning strategies to tackle complex learning issues under unreliable annotations effectively.
@@
From this perspective, our work presents a novel framework aimed at developing a transferable model adept at managing unreliable annotations, as illustrated in Figure 1. This model specifically targets the recognition of neurodegenerative disorders. These include neurocognitive disorders, represented by mild cognitive impairment(MCI) and Alzheimer’s Disease(AD), and movement disorders, exemplified by Parkinson’s Disease(PD). We selected these three diseases due to the potential progression relationship between AD and MCI [14] as well as the observed comorbidity between PD and cognitive impairment [15, 16]. The proposed framework is characterized by its innovative use of Manifold Attention and Confidence Stratification (MACS), which optimally synergizes modules including the Augmentor, Switcher, Encoder, Projector, and Stratifier. MACS’s success in achieving transferability across centers and diseases under unreliable annotations can be attributed to three key factors: 1) The establishment of an optimized mapping that leverages the strengths of both Euclidean and Riemannian manifold spaces enables the model to extract more effective EEG representations. 2) The integration of supervised and self-supervised learning strategies through confidence stratification to address the issue of learning from unreliable annotations. 3) An effective encoder enhances the representation space, which improves confidence stratification accuracy, thereby enabling the encoder to develop more robust and inductive representations in a mutually reinforcing cycle. Our primary contributions are outlined below:
@@ 2.1. Learning with Unreliable Annotations
Noisy label learning contains model-free and model-based strategies aimed at reducing the influence of incorrect labels. These strategies involve estimating noise patterns and conducting supervised learning with clean samples [17], as well as handling noisy labels and refining the model by addressing internal conflicts [18, 19]. Such methods depend on the analysis of the relationships within noisy data. Recent research in time-series [20] and image domains [21, 22] applies unique principles to delineate relationships. In the image domain, Promix [21] adopts small-loss criteria and prediction consistency to filter high-confidence examples, followed by learning through Debiased Semi-Supervised Training. Sel-CL [22] employs nearest neighbors to select confident pairs for supervised contrastive learning. In the time-series domain, CTW [20] selects confident examples based on small-loss criteria and applies time-warping to these instances to learn more robust representations. We use these to benchmark MACS and, following [23, 24], incorporate mix-up techniques [25, 26] to enhance learning from noisy labels.
@@ 2.2. Contrastive Learning for Time Series
Contrastive learning has demonstrated effectiveness in both unsupervised (e.g., SimCLR [27]) and supervised (e.g., SupCon [28]) contexts. It provides valuable self-learning strategies for pre-training in time series data, such as enhancing time-frequency [29] and cross-sample temporal consistency [3]. However, the impact of consistency pre-training on handling unreliable annotated data is less pronounced. Augmentation-based contrast [30] shows promise for semi-supervised learning with missing labels but is less effective without true label priors. To overcome these limitations, MACS introduces multi-view contrastive learning, guided by estimated confidence levels.
@@ 2.3. Manifold Learning for EEG Signals
Manifold-based modeling excels in EEG signal analysis for brain-computer interfaces (BCI) [31, 32], leveraging Riemannian geometry for high-dimensional neural data representation through affine-invariant geometric distances [33, 34]. Using EEG’s Riemannian structure for BCI domain-adaptation [10, 35, 36] and multi-task application [11] indicates the potential of Manifold geometry learning in identifying robust latent spaces. Such evidence bolsters the Encoder’s design in MACS. Our work advances current geometry learning by integrating dynamic functional networks with cross-attention mechanisms. Additionally, findings from our pilot study suggest that when integrated with contrastive learning, Riemannian geometry learning offers a promising alternative to GNNs [37].
@@ 3. MACS
In this section, we elucidate the methodology for learning from brain signals annotated with unreliable labels using the Manifold Attention and Confidence Stratification (MACS) framework. As depicted in Figure 2, MACS handles unreliable labels and enhances cross-center transferability through its modules (in Section 3.1), adaptive constraints (in Section 3.2), and training objective(in Section 3.3).
@@ 3.1. MACS Modules
In this module, data augmentation is applied to the preprocessed brain data as described in [29], following the equation:
@@
where $\bm{\epsilon}_{ij}\sim\mathcal{N}(0,\sigma^{2})$ is independently and identically distributed, with $\sigma$ representing the standard deviation. Additionally, $A_{d,T_{s}}$ is a segmented, non-overlapping fragment of $\bm{A}_{d,T}$, utilized to reduce computational load and increase the system’s robustness. Two random transformations are concurrently executed, resulting in dual brain variants. Differing from the method of specific weak and strong augmentations for two views [38], our approach operates a random configuration, leading to improved performance.
@@
In the Switcher module, dual brain variants undergo selective processing. As in Figure 3, distrusted inputs are directly forwarded to subsequent modules, while trusted ones are routed to a blender for generating interpolated samples. Inspired by the benefits of sample and network mixing, as discussed in [25] and [23], and aiming to enhance representation learning with partial labels and reduce confirmation bias in noisy data, we distort each trusted sample $A_{d,T_{s}}^{*}$ and its corresponding label $y^{*}$ by interpolating it with another randomly selected sample $(A_{d,T_{s}}^{+},y^{+})$, to produce blended variants, as described in Eq.(3). $\lambda$ is sampled from a uniform Beta(1, 1) distribution and adjusted to ensure it is always at least 0.5 by setting it to the greater value between $\lambda$ and $1-\lambda$.
@@
Differing from the referenced studies [25, 23], our method presents two distinctions: 1) We produce dual-blended brain variants while employing shared network modules, thus effectively decreasing model complexity. 2) We initiate data blending exclusively on samples exhibiting higher confidence levels, aiming to mitigate concerns related to unreliable labels. The confidence level associated with each sample in the mini-batch is evaluated using the Stratifier module in the MACS framework.
@@
The Encoder module $f_{En}$, grounded in our feature engineering findings (Section 4.3), initiates with a convolution starter $g_{str}$. This component reduces noise/artifacts and captures specific wave patterns (e.g., theta, alpha rhythms) and time-domain features of brain activity using convolutions and normalization, effectively decoding brain signals at a lower level[39]. This process yields $X_{d,T_{s}}=g_{\text{str}}(A_{d,T_{s})}$. Subsequently, a temporal clipper $g_{\text{clp}}$ is applied to these low-level features $X_{d,T_{s}}$ across the sampling dimension. This operation enhances their representation by leveraging the high temporal resolution of EEG signals.A series of feature clips $\{\ X_{dim_{1},t_{i}}\}\ _{i=1}^{I}$ are then processed by a synchronization analyzer $g_{syn}$, which measures the coupling between pairs of $d$ sensors, ultimately producing a dynamic functional network represented as $\Phi=\left[\phi_{t_{1}},\ldots,\phi_{t_{i}},\ldots,\phi_{t_{I}}\right]$.
With evidence from learning Riemannian geometry for structured brain data as shown in [36, 11], a manifold converter $g_{mfd}$ transforms dynamic network $\Phi$ into a set of symmetric positive definite (SPD) matrices, $S_{d}^{++}(\mathbb{R})$, thus defining a Riemannian manifold $\mathcal{M}$. This transformation involves eigenvalue decomposition and transpose manipulation of each adjacency matrix, represented as $\mathcal{D}=g_{mfd}(\Phi)$. To measure those $S_{d}^{++}$, we utilized the Log-Euclidean Riemannian metric [33], as it serves as the first-order approximation of the Affine-Invariant metric. At each point $\mathcal{D}_{t_{i}}$ on the manifold $\mathcal{M}$, a tangent space is defined by $\log:\mathcal{M}\to\mathcal{T}_{\mathcal{M}}$, where the inverse is operated by exponential mapping, as illustrated in Figure 4. The distance between any two points on the manifold $\mathcal{M}$ is then calculated using the following equation:
@@
Utilizing this metric to reflect the correlation between points $\mathcal{D}_{t_{i}}\in S_{d}^{++}$, we constructed a manifold-based dynamic attention block $g_{datt}$, following the approach in [11], but with modifications to incorporate a cross-temporal attention mechanism. This integration of dynamic relationships is mathematically represented as:
@@
where $\delta$ scales the values to a range of 0 to 1. $f_{W_{k}}$ represents the operation of $W_{k}$ multiplied by a matrix and then by $W_{k}^{\top}$. This can be expressed as $f_{W_{k}}(X)=W_{k}XW_{k}^{\top}$. Similarly, $f_{W_{q}}$ and $f_{W_{v}}$ operate in the same manner, where $W_{q}$, $W_{k}$, and $W_{v}\in\mathbb{R}^{d\times d_{1}}$ are learned weight matrices used for the bilinear mapping. Subsequently, the fused feature matrices =$\mathcal{F}=\{\mathcal{F}_{i}\}$ are subjected to a nonlinear activation process, involving the rectification of eigenvalues, as detailed in [33]. To enable the measurement of these features $\mathcal{F}$ in Euclidean space for further processing, the embeddings $\hat{\mathcal{F}}$ are obtained by mapping $\mathcal{F}$ onto the tangent space, followed by flattening and concatenating them together.
@@
  • Considering a sample trusted if its neighbor-determined label matches its original label. Following [40], we also adopted a dynamic thresholding approach to ensure class balance.

### v2

@@ Abstract.
Cross-center data heterogeneity and annotation unreliability significantly challenge the intelligent diagnosis of diseases using brain signals. A notable example is the EEG-based diagnosis of neurodegenerative diseases, which features subtler abnormal neural dynamics typically observed in small-group settings. To advance this area, in this work, we introduce a transferable framework employing Manifold Attention and Confidence Stratification (MACS) to diagnose neurodegenerative disorders based on EEG signals sourced from four centers with unreliable annotations. The MACS framework’s effectiveness stems from these features: 1) The Augmentor generates various EEG-represented brain variants to enrich the data space; 2) The Switcher enhances the feature space for trusted samples and reduces overfitting on incorrectly labeled samples; 3) The Encoder uses the Riemannian manifold and Euclidean metrics to capture spatiotemporal variations and dynamic synchronization in EEG; 4) The Projector, equipped with dual heads, monitors consistency across multiple brain variants and ensures diagnostic accuracy; 5) The Stratifier adaptively stratifies learned samples by confidence levels throughout the training process; 6) Forward and backpropagation in MACS are constrained by confidence stratification to stabilize the learning system amid unreliable annotations. Our subject-independent experiments, conducted on both neurocognitive and movement disorders using cross-center corpora, have demonstrated superior performance compared to existing related algorithms. This work not only improves EEG-based diagnostics for cross-center and small-setting brain diseases but also offers insights into extending MACS techniques to other data analyses, tackling data heterogeneity and annotation unreliability in multimedia and multimodal content understanding. We have released our code here: https://github.com/ICI-BCI/EEG-Disease-MACS.
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3681645isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Modeling methodologiesccs: Applied computing Health care information systems
@@ 1. Introduction
Addressing the challenges posed by data heterogeneity and annotation unreliability is essential for the analysis and interpretation of multimedia and multimodal data (Wang et al., 2023b; Zhang et al., 2023b). These issues are particularly prevalent in human-centered computing, where they encounter even greater difficulties in small-group settings. A prime example is the analysis of neural dynamics, such as the automated diagnosis of cognitive and movement brain disorders through electroencephalogram (EEG) signals  (Wang et al., 2023a; Chang et al., 2023). While EEG-based diagnosis models have undergone extensive research, they are often disease-specific, rely on single-center studies, and depend on fully supervised learning with annotated data  (Yasin et al., 2021). Developing effective and transferable models capable of handling cross-center data and low-quality annotations is crucial for enhancing adaptability across different diseases and advancing EEG modeling techniques.
@@
Current EEG analyzing methods range from traditional feature engineering to deep learning approaches. Various features extracted from domains like time (e.g., Hjorth parameters), frequency (e.g., Fourier Transform), dynamics (e.g., entropy), and functional networks (e.g., synchronization and phase coupling) have demonstrated efficacy due to EEG’s intrinsic characteristics (Coelho et al., 2023; Cavanagh et al., 2018; Suuronen et al., 2023; Heinrichs et al., 2023). The sparsity of handcrafted features has led to the adoption of deep learning models, such as convolutional neural networks, graph neural networks, and hybrid approaches, to capture spatial-temporal patterns, structured characteristics, and other high-level representations (Ju and Guan, 2023; Pan et al., 2022; Cai et al., 2023; Zhang et al., 2023a). However, these existing methods often struggle with noisy data and lack cross-center applicability. This may primarily be due to: 1) the ineffective mapping that cannot project EEG signals into a high-level feature space invariant across diverse data distributions from multiple centers; 2) the absence of integrating advanced learning strategies to tackle complex learning issues under unreliable annotations effectively.
@@
From this perspective, our work presents a novel framework aimed at developing a transferable model adept at managing unreliable annotations, as illustrated in Figure 1. This model specifically targets the recognition of neurodegenerative disorders. These include neurocognitive disorders, represented by mild cognitive impairment(MCI) and Alzheimer’s Disease(AD), and movement disorders, exemplified by Parkinson’s Disease(PD). We selected these three diseases due to the potential progression relationship between AD and MCI (Knopman et al., 2021) as well as the observed comorbidity between PD and cognitive impairment (Dag et al., 2021; Anjum et al., 2024). The proposed framework is characterized by its innovative use of Manifold Attention and Confidence Stratification (MACS), which optimally synergizes modules including the Augmentor, Switcher, Encoder, Projector, and Stratifier. MACS’s success in achieving transferability across centers and diseases under unreliable annotations can be attributed to three key factors: 1) The establishment of an optimized mapping that leverages the strengths of both Euclidean and Riemannian manifold spaces enables the model to extract more effective EEG representations. 2) The integration of supervised and self-supervised learning strategies through confidence stratification to address the issue of learning from unreliable annotations. 3) An effective encoder enhances the representation space, which improves confidence stratification accuracy, thereby enabling the encoder to develop more robust and inductive representations in a mutually reinforcing cycle. Our primary contributions are outlined below:
@@ 2.1. Learning with Unreliable Annotations
Noisy label learning contains model-free and model-based strategies for reducing the influence of incorrect labels. These strategies involve estimating noise patterns and conducting supervised learning with clean samples (Hendrycks et al., 2018), as well as handling noisy labels and refining the model by addressing internal conflicts (Malach and Shalev-Shwartz, 2017; Tan et al., 2021). Such methods depend on the analysis of the relationships within noisy data. Recent research in time-series (Ma et al., 2023) and image domains (Xiao et al., 2023; Li et al., 2022) applies unique principles to delineate relationships. In the image domain, Promix (Xiao et al., 2023) adopts small-loss criteria and prediction consistency to filter high-confidence examples, followed by learning through Debiased Semi-Supervised Training. Sel-CL (Li et al., 2022) employs nearest neighbors to select confident pairs for supervised contrastive learning. In the time-series domain, CTW (Ma et al., 2023) selects confident examples based on small-loss criteria and applies time-warping to these instances to learn more robust representations. We use these to benchmark MACS and, following (Li et al., 2020; Jiang et al., 2020), incorporate mix-up techniques (Berthelot et al., 2019; Zhang et al., 2021) to enhance learning from noisy labels.
@@ 2.2. Contrastive Learning for Time Series
Contrastive learning has demonstrated effectiveness in both unsupervised (e.g., SimCLR (Chen et al., 2020)) and supervised (e.g., SupCon (Khosla et al., 2020)) contexts. It provides valuable self-learning strategies for pre-training in time series data, such as enhancing time-frequency (Zhang et al., 2022a) and cross-sample temporal consistency (Wang et al., 2023a). However, the impact of such pre-training on handling unreliable annotated data is less pronounced. Augmentation-based contrast (Eldele et al., 2023) shows promise for semi-supervised learning with missing labels but is less effective without true label priors. Therefore, our proposed MACS introduces multi-view contrastive learning to overcome these limitations, guided by estimated confidence levels.
@@ 2.3. Manifold Learning for EEG Signals
Manifold-based modeling excels in EEG signal analysis for brain-computer interfaces (BCI) (Yger et al., 2016; Chevallier et al., 2021), leveraging Riemannian geometry for high-dimensional neural data representation through affine-invariant geometric distances (Huang and Van Gool, 2017; Sabbagh et al., 2019). Using EEG’s Riemannian structure for BCI domain-adaptation (Ju and Guan, 2023; Kobler et al., 2022; Bonet et al., 2023) and multi-task application (Pan et al., 2022) indicates the potential of Manifold geometry learning in identifying robust latent spaces. Such evidence bolsters the Encoder’s design in MACS. Our work advances current geometry learning by integrating dynamic functional networks with cross-attention mechanisms. Additionally, findings from our pilot study suggest that when integrated with contrastive learning, Riemannian geometry learning offers a promising alternative to GNNs (Zhang et al., 2022b).
@@ 3. MACS
In this section, we elucidate the methodology for learning from brain signals annotated with unreliable labels using the Manifold Attention and Confidence Stratification (MACS) framework. As depicted in Figure 2, MACS handles unreliable labels and enhances cross-center transferability through its modules (in Sec. 3.1), adaptive constraints (in Sec. 3.2), and training objective(in Sec. 3.3).
@@ 3.1. MACS Modules
In this module, data augmentation is applied to the preprocessed brain data as described in (Zhang et al., 2022a), following the equation:
@@
where $\bm{\epsilon}_{ij}\sim\mathcal{N}(0,\sigma^{2})$ is independently and identically distributed, with $\sigma$ representing the standard deviation. Additionally, $A_{d,T_{s}}$ is a segmented, non-overlapping fragment of $\bm{A}_{d,T}$, utilized to reduce computational load and increase the system’s robustness. Two random transformations are concurrently executed, resulting in dual brain variants. Differing from the method of specific weak and strong augmentations for two views (Eldele et al., 2021), our approach operates a random configuration, leading to improved performance.
@@
In the Switcher module, dual brain variants undergo selective processing. As in Figure 3, distrusted inputs are directly forwarded to subsequent modules, while trusted ones are routed to a blender for generating interpolated samples. Inspired by the benefits of sample and network mixing, as discussed in (Berthelot et al., 2019) and (Li et al., 2020), and aiming to enhance representation learning with partial labels and reduce confirmation bias in noisy data, we distort each trusted sample $A_{d,T_{s}}^{*}$ and its corresponding label $y^{*}$ by interpolating it with another randomly selected sample $(A_{d,T_{s}}^{+},y^{+})$, to produce blended variants, as described in Eq.(3). $\lambda$ is sampled from a uniform Beta(1, 1) distribution and adjusted to ensure it is always at least 0.5 by setting it to the greater value between $\lambda$ and $1-\lambda$.
@@
Differing from the referenced studies (Berthelot et al., 2019; Li et al., 2020), our method presents two distinctions: 1) We produce dual-blended brain variants while employing shared network modules, thus effectively decreasing model complexity. 2) We initiate data blending exclusively on samples exhibiting higher confidence levels, aiming to mitigate concerns related to unreliable labels. The confidence level associated with each sample in the mini-batch is evaluated using the Stratifier module in the MACS framework.
@@
The Encoder module $f_{En}$, grounded in our feature engineering findings (Section 4.3), initiates with a convolution starter $g_{str}$. This component reduces noise/artifacts and captures specific wave patterns (e.g., theta, alpha rhythms) and time-domain features of brain activity using convolutions and normalization, effectively decoding brain signals at a lower level(Défossez et al., 2023). This process yields $X_{d,T_{s}}=g_{\text{str}}(A_{d,T_{s})}$. Subsequently, a temporal clipper $g_{\text{clp}}$ is applied to these low-level features $X_{d,T_{s}}$ across the sampling dimension. This operation enhances their representation by leveraging the high temporal resolution of EEG signals.A series of feature clips $\{\ X_{dim_{1},t_{i}}\}\ _{i=1}^{I}$ are then processed by a synchronization analyzer $g_{syn}$, which measures the coupling between pairs of $d$ sensors, ultimately producing a dynamic functional network represented as $\Phi=\left[\phi_{t_{1}},\ldots,\phi_{t_{i}},\ldots,\phi_{t_{I}}\right]$.
With evidence from learning Riemannian geometry for structured brain data as shown in (Bonet et al., 2023; Pan et al., 2022), a manifold converter $g_{mfd}$ transforms dynamic network $\Phi$ into a set of symmetric positive definite (SPD) matrices, $S_{d}^{++}(\mathbb{R})$, thus defining a Riemannian manifold $\mathcal{M}$. This transformation involves eigenvalue decomposition and transpose manipulation of each adjacency matrix, represented as $\mathcal{D}=g_{mfd}(\Phi)$. To measure those $S_{d}^{++}$, we utilized the Log-Euclidean Riemannian metric (Huang and Van Gool, 2017), as it serves as the first-order approximation of the Affine-Invariant metric. At each point $\mathcal{D}_{t_{i}}$ on the manifold $\mathcal{M}$, a tangent space is defined by $\log:\mathcal{M}\to\mathcal{T}_{\mathcal{M}}$, where the inverse is operated by exponential mapping, as illustrated in Figure 4. The distance between any two points on the manifold $\mathcal{M}$ is then calculated using the following equation:
@@
Utilizing this metric to reflect the correlation between points $\mathcal{D}_{t_{i}}\in S_{d}^{++}$, we constructed a manifold-based dynamic attention block $g_{datt}$, following the approach in (Pan et al., 2022), but with modifications to incorporate a cross-temporal attention mechanism. This integration of dynamic relationships is mathematically represented as:
@@
where $\delta$ scales the values to a range of 0 to 1. $f_{W_{k}}$ represents the operation of $W_{k}$ multiplied by a matrix and then by $W_{k}^{\top}$. This can be expressed as $f_{W_{k}}(X)=W_{k}XW_{k}^{\top}$. Similarly, $f_{W_{q}}$ and $f_{W_{v}}$ operate in the same manner, where $W_{q}$, $W_{k}$, and $W_{v}\in\mathbb{R}^{d\times d_{1}}$ are learned weight matrices used for the bilinear mapping. Subsequently, the fused feature matrices =$\mathcal{F}=\{\mathcal{F}_{i}\}$ are subjected to a nonlinear activation process, involving the rectification of eigenvalues, as detailed in (Huang and Van Gool, 2017). To enable the measurement of these features $\mathcal{F}$ in Euclidean space for further processing, the embeddings $\hat{\mathcal{F}}$ are obtained by mapping $\mathcal{F}$ onto the tangent space, followed by flattening and concatenating them together.
@@
  • Considering a sample trusted if its neighbor-determined label matches its original label. Following (Ortego et al., 2021), we also adopted a dynamic thresholding approach to ensure class balance.